7.2 Draw the major organic product(s) of the following reaction. NaNH / NH3(I) CH3CH2CH2CH2-Br CH3CH2-CEC-H • You do not have to

b) Draw the organic product(s) of the following reaction. NaNH2 NH (1) -csc-H C-H • You do not have to consider stereochemistry.

c)Draw the major organic product(s) of the following reaction. Br met ee CH3 H3C-C-CEC-H het, som NaNH, NH (1) CH3CH2-CH-CH2 va
Answer #1

7.2 P CHE-CH-C=C-H NaNH₂ ? CH4-C1-C=C Na tots tenet Bo CH₂=CH-C=C- CH2 - CH - CH2 - CHA Terminal alkynes bearing a Hy doogen